A bridge is supported by two piers located 12 meters apart. Both the left and right piers provide an upward force on the bridge, labeled FL and FR respectively.

a. If a 1700 kg car comes to rest at a point 5 meters from the left pier, how much force (in N) will the bridge provide to the left and right piers? (Enter the magnitudes.)

FL = N

FR =   N

How will FL and FR change as the car drives to the right side of the bridge?

A. As the car moves to the right side of the bridge, FL will increase and FR will decrease.

B. As the car moves to the right side of the bridge, both FL and FR will decrease.     

C. As the car moves to the right side of the bridge, both FL and FR will increase.

D. As the car moves to the right side of the bridge, FL will decrease and FR will increase.
Answer #1

distance apart the piers = 12 m

weight of the car = 1700g = 16660 N

distance the car from left = 5 m

taking moments about left pier

M = FR *12 - 16660 *5 =0

FR = 6942 N

taking moments about the right pier

M = FL *12 - 16660 *(12-5) =0

FL = 9718 N

b) (D) - As the car moves to the FR increases and FL decreases.

distance of the load from the right decrease hence Fr increases